# Detector Response to Gravitational Wave Polarizations in Gravitational Quantum Field Theory

### v1

@@ 2.3 Response on the detectors
While the interference response typically measures differential changes between adjacent arms, our analysis primarily focuses on the distinctive breathing mode in GQFT, which is a transverse and isotropic polarization. Since interference subtraction could potentially suppress this mode’s signature, we instead examine single-arm responses through the relative length variation $\dfrac{\delta\ell_{ij}}{\ell_{ij}}$. This approach, combined with a metric perturbation expansion by polarization components, enables isolated examination of each polarization’s contribution.
For clarity and conciseness, we restrict our analysis to periodic detector effects and GWs propagating along the SSB frame’s z-direction. When the GW coordinate system aligns with the detector frame, the arm responses to polarization modes simplify considerably. Let us consider the GW source as a monochromatic signal, with the amplitude response for three polarization modes (generated by compact binary systems like double white dwarfs) given by the following expressions Jin & Qiao (2024),
$A:=\left(\frac{m_{1}m_{2}}{\left(m_{1}+m_{2}\right)^{2}}\right)^{3/5}\left(m_{1}+m_{2}\right)\omega.$ (18)
The detector arm responses to different polarization modes, $h_{ij,+}(t),h_{ij,\times}(t)$ and $h_{ij,b}(t)$, are given by:
$h_{12,+}(t)=\frac{A^{5/3}}{576\omega d_{L}}\Bigg{(}3\left(2\sin\left(\pi\left(\frac{1}{6}-4t\right)\right)+2\cos(4\pi t)-9\right)^{2}\left(\cos^{2}(\theta)\cos^{2}(\varphi)-\sin^{2}(\varphi)\right)+\\ \left(2\sqrt{3}\sin(4\pi t)+2\sqrt{3}\cos\left(\pi\left(\frac{1}{6}-4t\right)\right)+9\right)^{2}\left(\cos^{2}(\theta)\sin^{2}(\varphi)-\cos^{2}(\varphi)\right)+12\sqrt{3}\sin(2\theta)\\ \left(\cos(2\pi t)+\cos\left(\frac{1}{3}\pi(6t+1)\right)\right)\cos(\varphi)\left(2\sin\left(\pi\left(\frac{1}{6}-4t\right)\right)+2\cos(4\pi t)-9\right)+\frac{3}{2}(\cos(2\theta)+3)\sin(2\varphi)\\ \left(2\sin(4\pi t)+2\cos\left(\pi\left(\frac{1}{6}-4t\right)\right)+3\sqrt{3}\right)\left(2\sin\left(\pi\left(\frac{1}{6}-4t\right)\right)+2\cos(4\pi t)-9\right)+12\sin(2\theta)\\ \left(\cos(2\pi t)+\cos\left(\frac{1}{3}\pi(6t+1)\right)\right)\sin(\varphi)\left(2\sqrt{3}\sin(4\pi t)+2\sqrt{3}\cos\left(\pi\left(\frac{1}{6}-4t\right)\right)+9\right)\\ +144\sin^{2}(\theta)\left(\cos(2\pi t)+\cos\left(\frac{1}{3}\pi(6t+1)\right)\right)^{2}\Bigg{)},$ (19)
$h_{12,\times}(t)=\frac{A^{5/3}}{48\omega d_{L}}\Bigg{(}4\sin(\theta)\left(\cos(2\pi t)+\cos\left(\frac{1}{3}\pi(6t+1)\right)\right)\\ \left(\sqrt{3}\left(2\left(\sin(4\pi t-\varphi)+\cos\left(-4\pi t+\varphi+\frac{\pi}{6}\right)\right)+9\sin(\varphi)\right)+9\cos(\varphi)\right)\\ +3\cos(\theta)\left(-12\sin(4\pi t-2\varphi)+\sin(8\pi t-2\varphi)+\sqrt{3}\cos(8\pi t-2\varphi)-9\sin(2\varphi)-9\sqrt{3}\cos(2\varphi)\right)\Bigg{)},$ (20)
and
$h_{12,b}(t)=\frac{A^{4/3}}{1152\omega d_{L}}\Bigg{(}9\sin^{2}(\theta)\cos(2\varphi)\left(\sqrt{3}\sin(8\pi t)+12\cos(4\pi t)-\cos(8\pi t)-9\right)-9\sin^{2}(\theta)\sin(2\varphi)\\ \left(-12\sin(4\pi t)+\sin(8\pi t)+\sqrt{3}\cos(8\pi t)\right)+9\sin^{2}(\theta)\left(16\left(\cos(2\pi t)+\cos\left(\frac{1}{3}\pi(6t+1)\right)\right)^{2}+9\sqrt{3}\sin(2\varphi)\right)\\ +12\sqrt{3}\sin(2\theta)\left(\cos(2\pi t)+\cos\left(\frac{1}{3}\pi(6t+1)\right)\right)\cos(\varphi)\left(2\sin\left(\pi\left(\frac{1}{6}-4t\right)\right)+2\cos(4\pi t)-9\right)\\ +12\sin(2\theta)\left(\cos(2\pi t)+\cos\left(\frac{1}{3}\pi(6t+1)\right)\right)\sin(\varphi)\left(2\sqrt{3}\sin(4\pi t)+2\sqrt{3}\cos\left(\pi\left(\frac{1}{6}-4t\right)\right)+9\right)\\ +9(\cos(2\theta)+3)\left(-3\sin\left(\pi\left(\frac{1}{6}-4t\right)\right)+\sqrt{3}\sin(4\pi t)+\sqrt{3}\cos\left(\pi\left(\frac{1}{6}-4t\right)\right)-3\cos(4\pi t)+10\right)\Bigg{)},$ (21)
for the detector Arm12, and
$h_{13,+}(t)=\frac{A^{5/3}}{1152\omega d_{L}}\Bigg{(}9(\cos(2\theta)+3)\cos(2\varphi)\left(\sqrt{3}\sin(8\pi t)-12\cos(4\pi t)+\cos(8\pi t)+9\right)+24\sin(2\theta)\\ \left(\sin\left(\pi\left(2t+\frac{1}{6}\right)\right)+\cos(2\pi t)\right)\left(2\sqrt{3}\sin\left(4\pi t-\varphi+\frac{\pi}{6}\right)+2\sqrt{3}\cos(4\pi t-\varphi)-9\sin(\varphi)-9\sqrt{3}\cos(\varphi)\right)\\ -9(\cos(2\theta)+3)\sin(2\varphi)\left(12\sin(4\pi t)-\sin(8\pi t)+\sqrt{3}\cos(8\pi t)-9\sqrt{3}\right)\\ +36\sin^{2}(\theta)\left(-4\sin\left(\pi\left(\frac{1}{6}-4t\right)\right)+7\sqrt{3}\sin(4\pi t)+11\cos(4\pi t)+2\right)\Bigg{)},$ (22)
$h_{13,\times}(t)=-\frac{A^{5/3}}{48\omega d_{L}}\Bigg{(}4\sin(\theta)\left(\sin\left(\pi\left(2t+\frac{1}{6}\right)\right)+\cos(2\pi t)\right)\\ \left(\sqrt{3}(-2\sin(4\pi t-\varphi)-9\sin(\varphi))+2\sqrt{3}\cos\left(4\pi t-\varphi+\frac{\pi}{6}\right)+9\cos(\varphi)\right)\\ +3\cos(\theta)\left(12\sin(4\pi t-2\varphi)-\sin(8\pi t-2\varphi)+\sqrt{3}\cos(8\pi t-2\varphi)+9\sin(2\varphi)-9\sqrt{3}\cos(2\varphi)\right)\Bigg{)},$ (23)
and
$h_{13,b}(t)=\frac{A^{4/3}}{1152\omega d_{L}}\Bigg{(}\left(-2\sqrt{3}\sin(4\pi t)+2\sqrt{3}\cos\left(\pi\left(4t+\frac{1}{6}\right)\right)+9\right)^{2}\left(\cos^{2}(\theta)\sin^{2}(\varphi)+\cos^{2}(\varphi)\right)\\ +3\left(2\sin\left(\pi\left(4t+\frac{1}{6}\right)\right)+2\cos(4\pi t)-9\right)^{2}\left(\cos^{2}(\theta)\cos^{2}(\varphi)+\sin^{2}(\varphi)\right)\\ +3\sin^{2}(\theta)\sin(2\varphi)\left(-2\sin(4\pi t)+2\cos\left(\pi\left(4t+\frac{1}{6}\right)\right)+3\sqrt{3}\right)\left(2\sin\left(\pi\left(4t+\frac{1}{6}\right)\right)+2\cos(4\pi t)-9\right)\\ -12\sin(2\theta)\sin(\varphi)\left(\sin\left(\pi\left(2t+\frac{1}{6}\right)\right)+\cos(2\pi t)\right)\left(-2\sqrt{3}\sin(4\pi t)+2\sqrt{3}\cos\left(\pi\left(4t+\frac{1}{6}\right)\right)+9\right)\\ +12\sqrt{3}\sin(2\theta)\cos(\varphi)\left(\sin\left(\pi\left(2t+\frac{1}{6}\right)\right)+\cos(2\pi t)\right)\left(2\sin\left(\pi\left(4t+\frac{1}{6}\right)\right)+2\cos(4\pi t)-9\right)\\ +144\sin^{2}(\theta)\left(\sin\left(\pi\left(2t+\frac{1}{6}\right)\right)+\cos(2\pi t)\right)^{2}\Bigg{)},$ (24)
for the detector Arm13, and
$h_{23,+}(t)=\frac{A^{5/3}}{48\omega d_{L}}\Bigg{(}-3\sqrt{3}\sin(2\theta)(5\sin(2\pi t)+\sin(6\pi t))\sin(\varphi)+\left(\cos\left(\pi\left(\frac{1}{6}-4t\right)\right)+\cos\left(\pi\left(4t+\frac{1}{6}\right)\right)+3\sqrt{3}\right)^{2}\\ \left(\cos^{2}(\theta)\sin^{2}(\varphi)-\cos^{2}(\varphi)\right)+3\sin^{2}(4\pi t)\left(\cos^{2}(\theta)\cos^{2}(\varphi)-\sin^{2}(\varphi)\right)\\ -\frac{1}{2}\sqrt{3}(\cos(2\theta)+3)\sin(4\pi t)\left(\cos\left(\pi\left(\frac{1}{6}-4t\right)\right)+\cos\left(\pi\left(4t+\frac{1}{6}\right)\right)+3\sqrt{3}\right)\sin(2\varphi)\\ +12\sqrt{3}\sin(\theta)\cos(\theta)\sin(2\pi t)\sin(4\pi t)\cos(\varphi)+36\sin^{2}(\theta)\sin^{2}(2\pi t)\Bigg{)}$ (25)
$h_{23,\times}(t)=-\frac{A^{5/3}}{8\omega d_{L}}\Bigg{(}4\sqrt{3}\sin(\theta)\sin(2\pi t)(\cos(4\pi t-\varphi)+3\cos(\varphi))+\cos(\theta)(6\sin(4\pi t-2\varphi)+\sin(8\pi t-2\varphi)-9\sin(2\varphi))\Bigg{)},$ (26)
and
$h_{23,b}(t)=\frac{A^{4/3}}{96\omega d_{L}}\Bigg{(}-3\sqrt{3}\sin(2\theta)(5\sin(2\pi t)+\sin(6\pi t))\sin(\varphi)\\ +\left(\cos\left(\pi\left(\frac{1}{6}-4t\right)\right)+\cos\left(\pi\left(4t+\frac{1}{6}\right)\right)+3\sqrt{3}\right)^{2}\left(\cos^{2}(\theta)\sin^{2}(\varphi)+\cos^{2}(\varphi)\right)\\ +3\sin^{2}(4\pi t)\left(\cos^{2}(\theta)\cos^{2}(\varphi)+\sin^{2}(\varphi)\right)\\ +\sqrt{3}\sin^{2}(\theta)\sin(4\pi t)\left(\cos\left(\pi\left(\frac{1}{6}-4t\right)\right)+\cos\left(\pi\left(4t+\frac{1}{6}\right)\right)+3\sqrt{3}\right)\sin(2\varphi)\\ +12\sqrt{3}\sin(\theta)\cos(\theta)\sin(2\pi t)\sin(4\pi t)\cos(\varphi)+36\sin^{2}(\theta)\sin^{2}(2\pi t)\Bigg{)},$ (27)
for the detector Arm23.
In the above formalisms, we neglect polarization mode mixing by setting $\psi=0$. The characteristic amplitude parameter is defined as $A:=\left(\frac{\text{m1}\text{m2}}{(\text{m1}+\text{m2})^{2}}\right)^{3/5}(\text{m1}+\text{m2})\omega$. Here, $m1$ and $m2$ are the component masses of the binary system, and $\omega$ is the GW frequency. Furthermore, $d_{L}$ denotes the luminosity distance, and $\theta$ and $\phi$ represent the angular sky coordinates of the source (polar and azimuthal angles, respectively).
3 Results
@@
We examine the specific case of HM Cancri (RX J0806.3+1527), with the following parameters: sky location: $(\theta,\varphi)=(1.65,2.10)$ radians, component masses: $0.55\mathrm{M}_{\odot}$ and $0.27\mathrm{M}_{\odot}$, GW frequency: $6.22\mathrm{mHz}$, and luminosity distance: $5\mathrm{kpc}$.
@@
The detector response depends on the relative orientation between the detector arms and the GW source, producing both positive and negative values. We therefore consider the absolute value of the relative length change $\delta l/l$ to characterize the signal. This allows us to separately analyze the responses to plus, cross, and breathing polarization modes over a one-year observation period, and to quantify the differences between GR and GQFT predictions.
@@
Figure 1: GW strain response of arm12 to the HM Cancri system at a specific sky location. Left panel: Relative strain components for plus, cross, and breathing polarization modes. Middle panel: Total strain comparison between GR (sum of plus and cross modes) and GQFT (including all three polarizations). Right panel: Temporal evolution over one year (normalized to 0-1), showing GR null points versus GQFT maxima for the source position in the sky map.
@@
Figure 3: Gravitational wave response of arm23, following the same presentation as Figure 1.
[…]
This null-point method offers a powerful alternative to direct mode separation, which is observationally challenging. The temporal separation between theory predictions (up to one month in some cases) provides a more robust test than amplitude-based discrimination. Our analysis focuses specifically on distinguishing GR and GQFT through polarization responses, making the wave source details largely irrelevant. However, the annual distribution of null points does encode source information for future study. For far-field GWs, only the polarization response matters.
For every sky location, there exist characteristic time periods when GR predicts null responses while the $b$ mode remains detectable as predicted by GQFT. These optimal observation times could be mapped across the sky, with colors representing the fractional year times of GR nulls and maximal GQFT signals.
4 Conclusion
@@
We model detector orbits and gravitational wave propagation using first-order approximations within the Solar System Barycenter (SSB) coordinate system. Through the parameterized post-Einsteinian (ppE) formalism, we derive explicit expressions for polarization amplitudes, yielding compact analytical solutions for detector responses in both Taiji and LISA configurations. Using the ultra-compact binary HM Cancri as our benchmark system, we conduct a systematic comparison of polarization responses between GR and GQFT. Our investigation reveals several key findings: Polarization mode interference generates distinctive null points that exhibit marked differences between GR and GQFT predictions. These null points serve as robust discriminants between the two theories, circumventing the observational challenges associated with direct b-mode detection. In the HM Cancri system, we identify consistent temporal separations of approximately two weeks between GR and GQFT null points across both arm12 and arm23 configurations. Additionally, through sky mapping of optimal observation times, we demonstrate that arm23 exhibits a comparatively weaker dependence on source position relative to arm12, highlighting its potential advantages for certain observational scenarios.

### v2

@@ II.3 Response on the detectors
While the interference response typically measures differential changes between adjacent arms, our analysis primarily focuses on the distinctive breathing mode in GQFT, which is a transverse and isotropic polarization. Since interference subtraction could potentially suppress this mode’s signature, we instead examine single-arm responses through the length strain itself. This approach, combined with a metric perturbation expansion by polarization components, enables isolated examination of each polarization’s contribution.
Furthermore, we consider that Eq.(22) requires reconsideration. Substituting into Eq.(10) indicates that the length strain depends on contributions from both the waveform $A:=\int h(t)\mathop{}\!\mathrm{d}t$ and the response function $F:=\frac{1}{2}\hat{r}_{ij}^{a}(t)\otimes\hat{r}_{ij}^{b}(t)\varepsilon_{ab}$. Thus, the length strain should be a linear combination of the amplitude and the response function $\delta strain=\sum_{i\in polarizations}A_{i}F_{i}$. All model dependencies are incorporated within the waveform , while the response function depends solely on the detector configuration and position.
Therefore, by providing a model-independent response function, we can effectively characterize the response of different polarization modes in the Taiji orbit. Thereafter, any waveform model simply provides the amplitude, which is then linearly combined with the corresponding polarization response function to yield the length strain. To clearly illustrate the performance of each polarization mode, we neglect polarization mode mixing by setting $\psi=0$, at this time the analytical expression for the response function in any direction is provided in the appendixA.
Specifically, plots of the response $F$ for different detection arms are provided. Here we take HM Cancri (RX J0806.3+1527) as an example GW source, located at sky position $(\theta,\varphi)=(1.65,2.10)$ radians. At this point, the response functions for all polarisation modes on different arms can be plotted as following Figure 1.
Figure 1: Response function with sky position $(\theta,\varphi)=(1.65,2.10)$ radians on arm12, arm13 and arm23, where blue, orange, green, red, purple, and brown correspond to plus, cross, x, y, breathing, and longitudinal responses respectively.
The response function demonstrates the effect of the Taiji orbit on gravitational wave signals. By comparing the three distinct detector arms, it can be observed that the tensor modes exhibit a relatively stable periodic variation with respect to the detector, maintaining both positive and negative responses across all arms. This stability demonstrates the robust nature of the tensor modes within metric perturbations. In contrast, the vector modes show significant variations in their projections across different arms. Specifically, the y mode occasionally presents responses that are either entirely positive or negative across different arms. This indicates that the vector projection is highly sensitive to the detector’s position. Finally, the scalar modes are consistently positive and regular regardless of the detector arm, which corresponds to the inherent property of scalars being largely independent of the detector’s projection.
III Results in GQFT
@@
Let us consider the GW source as a nearly monochromatic chirp signal(generated by compact binary systems like double white dwarfs). The selected instance is HM Cancri (RX J0806.3+1527), with the following parameters: sky location: $(\theta,\varphi)=(1.65,2.10)$ radians, component masses: $0.55\mathrm{M}_{\odot}$ and $0.27\mathrm{M}_{\odot}$, GW frequency: $6.22\mathrm{mHz}$, and luminosity distance: $5\mathrm{kpc}$.
@@
The detector response depends on the relative orientation between the detector arm and the GW source, length changes can better describe the signal. This allows us to separately analyze the responses to plus, cross, and breathing polarization modes over a one-year observation period, and to quantify the differences between GR and GQFT predictions.
Given the vast disparity in timescales—the source’s characteristic period being on the order of seconds, while the response function’s period is on the order of years—the resulting signal in the time domain is highly oscillatory. We employ conventional time-frequency analysis methods for signal extraction. As the chirp signal in this context is nearly monochromatic, we can effectively analyze the distinctions between different gravitational theories by plotting a slice of the corresponding spectrogram.
@@
Figure 2: GW strain response of arm12 to the HM Cancri system at a specific sky location. Left figure: length strain components for plus, cross, and breathing polarization modes. Middle figure: Total strain comparison between GR (mix of plus and cross modes) and GQFT (including all three polarizations). Right figure: Illustrating the relative temporal rate of strain amplitude change predicted by GR and GQFT can serve as a supplementary reference for the middle figure.
@@
Figure 4: Gravitational wave response of arm23, following the same representation as Figure 2.
The distinctions observed in Figure 2-4 are attributed to the relative orientation with respect to the detector. Specifically, arm13 exhibits a notable difference compared to the other two detector arms. Along this arm, the plus and cross polarizations demonstrate a marked symmetry in their distribution, and the b mode fluctuations remain minimal. Consequently, this leads to a significant divergence in the amplitude predicted by GR and GQFT. This suggests that, for the specific event RX J0806.3+1527, arm13 offers enhanced capable for detecting the b mode, thereby providing a superior means of distinguishing between the predictions of GR and GQFT.
It is clear that signals corresponding to both GR and GQFT can be extracted using time-frequency analysis for any source on the sky location. And it can be observed that the detection of polarization modes exhibits strong dependence on the event position. This naturally leads to the question of whether certain sky locations are more conducive to differentiating between these two theories. To quantify this distinguishability, we consider a ratio of the maximum and minimum amplitudes predicted by GR and GQFT $Ratio=\dfrac{\mathrm{Max}[Amp]}{\mathrm{Min}[Amp]}$ with above analysis. This ratio is chosen specifically to mitigate the dependence on waveform, thereby characterizing a relative intensity. Meanwhile, the rate-of-change curve presented in the third part of Figure 2-4 effectively identifies the distribution of extremum, which aids in the calculation of the amplitude ratio. Then compute the subtraction between the ratio of GR and GQFT $Ratio_{GQFT}-Ratio_{GR}$ at each sky position. This value provides a reference for distinguishing different theories, and captures b-mode information. Based on this, we plotted a skymap covering the entire sky location as shown in Figure 5.
Figure 5: The subtraction of the maxima and minima ratio between GR and GQFT over the entire sky location, where the color represents the log discrepancy, redder coloring indicates a greater difference, while bluer coloring indicates a smaller difference, where the difference is exponential. Namely, indicates B mode detectable on arm12, arm13 and arm23 respectively.
Skymap5 reveals that only specific fixed sky locations can effectively differentiate between GR and GQFT. The diagram shows several figure-eight structures, which correspond to some special cases, when the probe arm is perpendicular to the wavefront, there is no response. Because both gravitational waves in GR and GQFT are transversely polarized. Furthermore, the region surrounding this structure tends to exhibit a greater magnitude (redder color), which corroborates the earlier finding that the scalar mode is minimally affected by the detector’s position, thereby accentuating the contribution of the b mode in this area. Additionally, the distinguishable sky positions are widely distributed for arm 23, suggesting that this specific arm orbit has a higher sensitivity to the b mode, potentially offering a more promising avenue for the detection of signatures predicted by GQFT.
Our analysis focus specifically on distinguishing GR and GQFT through polarization responses, making the wave source details largely irrelevant. For far-field GWs, only the polarization response matters. The GR and GQFT responses appear broadly similar because the $b,+,\text{and}\times$ modes are all transverse, with the $b$ mode exhibiting isotropic transverse breathing. This symmetry makes it nearly impossible to simultaneously null both the + and $\text${ and $\times$ modes, motivating our focus on single-arm variations rather than interference measurements. Subtraction of simple interference of the dominant modes typically leaves only a small residual b-mode signal, suggesting that single-arm detection may be more sensitive to this polarization. In future, we shall consider more subtle interference configurations to enable better separation of the b mode.
IV Conclusion
@@
We model detector orbits and gravitational wave propagation using first-order approximations within the Solar System Barycenter (SSB) coordinate system. Through the parameterized post-Einsteinian (ppE) formalism, we derive explicit expressions for polarization amplitudes, yielding compact analytical solutions for detector responses in both Taiji configurations. We have derived the response function expression for any position and analysed its characteristics in tensor, vector and scalar modes , also using HM Cancri as an example in Figure 1. Using the ultra-compact binary HM Cancri as our benchmark system, we conduct a systematic comparison of polarization responses between GR and GQFT. Through time-frequency analysis of gravitational wave polarisation, we obtained the amplitudes on each detector arm under the Taiji orbit. And found that under the HM Cancri event, arm13 shows results distinct from the other arms, providing a methodology and evidence for detecting b mode in Figure 3. This also indicates that the observation of polarization modes is strongly correlated with the event localization. Additionally, through sky mapping of optimal observation event position, We demonstrate that only certain fixed positions can better separate different theories. Meanwhile, through skymap in Figure 5, we found that arm23 exhibits greater sensitivity on b mode, highlighting its potential advantages for certain observational scenarios.